The heating element of a flat iron provides a resistance of 24 ohm's when plugged into a 220-V outlet. Calculate the electric power consumed by the iron.


1
Expert's answer
2021-04-23T11:07:21-0400
"P=\frac{U^2}{R}\\P=\frac{220^2}{24}=2017\ W"